Can you help me to solve this three problems?

1. A total charge Q is distributed evenly over a thread of length l. The thread forms a semicircle. What is the potential in the center? (Consider that V = 0 at great distances).

2. It has a system with two loads struts of 25μC that are separated 6cm. The work that an external force must do to move a test load of 0.18μC from a point in the middle of the loads to another point that is 1cm closer to some of them is 0.5J.

3. Two identical point charges of 6.4 μC are initially 7.4 cm apart from one another. If they are released at the same time, how fast will they be moving when they are too far from one another? Suppose they have equal masses of 1.9 mg.
Answer #1

1) if radius of the semi circle is R

Then πR = l

So R = l/π

Now every elementary charge on thread is equi distant from the center of semi circle.

So potential at the center V = kQ/R = kQπ/l

Where k is electrostatic constant.